# Adaptive Smooth Non-Stationary Bandits

### v1

@@ 5   Gap-Dependent Dynamic Regret Bounds
where $L$ is the number of stationary phases, the $\ell$-th of which has gap profile $\{\delta_{\ell}(a)\}_{a\in[K]}$. Unfortunately, it’s long been known in the switching bandit literature that, in the worst case, Equation 3 cannot be attained simultaneously for different values of $L$ (Garivier and Moulines, 2011; Lattimore and Szepesvári, 2020). The reason for such a hardness is intuitively because of the additional exploration required to detect unknown changes, which forces $\sqrt{T}$ regret.
@@
In our next contribution, we give answers to both these questions in terms of the significant shift oracle, introduced in Section 4. Recalling such an oracle roughly plays arms until they incur significant regret Equation 1, and restarts at each significant shift, we’ll see that a careful regret analysis of this oracle gives rise to a faster rate than Equation 3 which is achievable adaptively in so-called safe environments (Definition 10).
@@ 5.1   Related Work on Gap-Dependent Regret
To start, we give an account of some works which aim to achieve the restarting oracle rate Equation 3 under structured non-stationarity:
@@
  • Seznec et al. (2020) achieve the rate Equation 3 in (restless) rotting bandits where rewards decrease in time.
@@ 5.2   Refined Regret Analysis of the Significant Shift Oracle
Notation 7.
Let $\mathcal{H}_{t}$ be be the $\sigma$-algebra generated by the random reward variables $\{Y_{s}(a)\}_{s\leq t,a\in[K]}$ and exogenous time-varying randomness $\{\pi_{s}\}_{s\leq t}$, as used by an algorithm $\pi$.
We’ll use $t_{1},\ldots,t_{K}$ to denote ordered stopping times with respect to the filtration $\{\mathcal{H}_{t}\}_{t\in[T]}$ and use $\mathcal{S}_{1},\cdots,\mathcal{S}_{T}$ to denote random subsets of $[K]$ which are adapted to this filtration.
@@
Let $t_{0}\doteq 1$ and let $\mathcal{S}_{1}\doteq[K]$. Then, we’ll recursively define $t_{i}$ and $\mathcal{S}_{t}$ for $t>t_{i-1}$ as follows: a stopping time $t_{i}>t_{i-1}$ is called an eviction time w.r.t. initial time $t_{i-1}$ if $\mathcal{S}_{t_{i}-1}=\mathcal{S}_{t_{i}-2}=\cdots=\mathcal{S}_{t_{i-1}}$ and
$\forall a\in\mathcal{S}_{t_{i}-1},[s_{1},s_{2}]\subseteq[1,t_{i}-1]:\sum_{s=s_{1}}^{s_{2}}\frac{\delta_{s}(a)}{|\mathcal{S}_{s}|}$ $\leq C_{2}\sqrt{\sum_{s=s_{1}}^{s_{2}}\frac{\log(T)}{|\mathcal{S}_{s}|}}.$ (4)
We call $(t_{1},t_{2},\ldots,t_{K})$ a sequence of eviction times with associated safe armsets $\mathcal{S}_{1}\supseteq\mathcal{S}_{2}\supseteq\cdots\supseteq\mathcal{S}_{T}$.
@@
Plainly speaking, Equation 5 captures the regret of the significant shift oracle, or specifically an elimination procedure which tracks the safe armsets $\mathcal{S}_{t}$ and uniformly explores $\mathcal{S}_{t}$ at round $t$. Note that Equation 5 is a random quantity as the $t_{i},\mathcal{S}_{t}$ may depend on the random rewards and exogenous randomness of some algorithm $\pi$. Despite this randomness, we’ll next show that for any valid $t_{i},\mathcal{S}_{t}$ satisfying Definition 8, the rate of Equation 5 recovers Equation 3 and is achievable adaptively in safe environments (Definition 10).
5.3   Properties of New Gap-Dependent Regret Rate
Proofs are deferred to Appendix C.
@@ 5.3   Properties of New Gap-Dependent Regret Rate
($\mathscr{R}(\{t_{i}\}_{i\in[K]},\{\mathcal{S}_{t}\}_{t\in[T]},\pi)\leq C_{2}^{2}\sum_{\ell=1}^{L}\sum_{a:\delta_{\ell}(a)>0}\frac{\log(T)}{\delta_{\ell}(a)}.$
We next show Equation 5 recovers the usual $\sqrt{KT}$ regret bound in safe environments. We first define such a notion.
@@
$\forall[s_{1},s_{2}]\subseteq[1,T]:\sum_{t=s_{1}}^{s_{2}}\frac{\delta_{t}({a^{\sharp}})}{|\mathcal{G}_{t}|}\leq C_{3}\sqrt{\sum_{t=s_{1}}^{s_{2}}\frac{\log(T)}{|\mathcal{G}_{t}|}}.$ (6)
@@
A significant shift cannot occur in a safe environment. Indeed, taking $\mathcal{G}_{t}\equiv[K]$ in Equation 6 gives us $\sum_{t=s_{1}}^{s_{2}}\delta_{t}({a^{\sharp}})<C_{3}\sqrt{K\cdot\log(T)\cdot(s_{2}-s_{1}+1)}$ which is a generalization of the reversal of Equation 1.
@@
($\mathscr{R}(\{t_{i}\}_{i\in[K]},\{\mathcal{S}_{t}\}_{t\in[T]},\pi)\leq C_{2}\sqrt{KT\log(T)}.$
@@ 5.4   Elimination Achieves Gap-Dependent Regret Rate in Safe Environments
As already hinted up to this point, we posit that a randomized variant of elimination (similar to that of Suk and Kpotufe, 2022) in fact attains the rate of Equation 5 in safe environments. We view our contribution here as not algorithmic, but rather a tighter regret analysis compared to Lemma 3 of Suk and Kpotufe (2022).
@@
We present (Algorithm 1) a slightly different version of Algorithm 2 of Suk and Kpotufe (2022). First, we estimate the relative gap $\delta_{t}(a^{\prime},a)$ via [^4] This estimator is slightly different than the importance-weighted one used in Suk and Kpotufe (2022); we require this form to mimic the form of Equation 4.:
$\hat{\delta}_{t}(a^{\prime},a)\doteq Y_{t}(a^{\prime})\cdot\boldsymbol{1}\{\pi_{t}=a^{\prime}\}-Y_{t}(a)\cdot\boldsymbol{1}\{\pi_{t}=a\}.$ (7)
Then, at round $t$, we’ll eliminate arms from an active armset $\mathcal{A}_{t}$ when an empirical analogue of Equation 4 based on $\hat{\delta}_{t}(a^{\prime},a)$ and $\mathcal{A}_{t}$ holds.
@@ 4       Evict bad arms:
5       $\mathcal{A}_{t}\leftarrow\mathcal{A}_{t}\backslash\left\{a\in[K]:\text{$\exists$ round $t_{0}\leq t$ s.t. $\underset{a^{\prime}\in\mathcal{A}_{t}}{\max}\sum_{s=t_{0}}^{t}\hat{\delta}_{s}(a^{\prime},a)>C_{5}\sqrt{\sum_{s=t_{0}}^{t}\frac{\log(T)}{|\mathcal{A}_{s}|}}$ holds}\right\}$.
@@
Given any safe bandit environment over $T$ rounds, letting $\pi$ be Algorithm 1, we have w.p. at least $1-1/T^{2}$, for some algorithm-dependent eviction times and safe armsets $\{t_{i}\}_{i\in[K]},\{\mathcal{S}_{t}\}_{t\in[T]}$:
$\sum_{t=1}^{T}\delta_{t}(\pi_{t})\leq C_{4}\cdot\left(\log(T)+\mathscr{R}(\{t_{i}\}_{i\in[K]},\{\mathcal{S}_{t}\}_{t\in[T]},\pi)\right).$
@@
On the other hand, as mentioned earlier, it’s known in switching bandits that the restarting oracle rate Equation 3 cannot be achieved adaptively for unknown $L$. However, this does not contradict our findings because the constructed hard environment (e.g. Lattimore and Szepesvári, 2020, Theorem 31.2) is not safe (Definition 10). Thus, similar to Suk and Kpotufe (2022), we find that the notion of significant shift (which decides the safeness of an environment) characterizes difficult non-stationarity in a new sense. So long as such a shift does not occur, we can attain the faster rate Equation 5.
@@ 5.5   Lower Bound for Gap-Dependent Regret Rate
We note the $\log(T)$ factor in Equation 4 of Definition 8 was only included for the sake of showing the regret upper bounds established up to this point. Going forward, we’ll ignore the $\log(T)$ factor when we refer to Equation 4.
@@
Let $\{t_{i}\}_{i=1}^{K}$ be an arbitrary set of rounds such that $t_{i+1}-t_{i}+1\geq K$ for all $i\in[K]$ with the convention that $t_{0}\doteq 1$ and $t_{K-1}=t_{K}\doteq T+1$. Fix a positive real number $R$ such that $R\leq\sum_{i=1}^{K-1}\sqrt{(t_{i}-t_{i-1})\cdot(K+1-i)}$. Let $\mathscr{E}$ be the class of environments such that (a) $t_{1},\ldots,t_{K}$ are valid deterministic eviction times with $C_{2}=(K-2)^{1/2}$ in Equation 4 for some shrinking sequence of safe armsets $\mathcal{S}_{1}\supseteq\cdots\supseteq\mathcal{S}_{T}$ and (b) such that:
@@ 6   Achievability of Gap-Dependent Rate in terms of Smoothness
Theorem 14 (Proof in Subsection C.5).
@@ Appendix A Proof of Dynamic Regret Lower Bound (Theorem 3)
The construction will rely on bump reward functions which also appear in the classical minimax lower bounds for integrated risk of nonparametric regression of $(\beta,L)$-Hölder functions (e.g., Section 2.5 of Tsybakov, 2009). This will be combined with a Le Cam’s method style of argument for establishing a regret lower bound for stationary bandits (Lattimore and Szepesvári, 2020, Theorem 15.2), which we’ll modify to work for segments of mild non-stationarity.
@@
At a high level, we’ll construct $M$ instances of a randomly selected $(\beta,\lambda)$-Hölder environment of length $T/M$. We’ll then argue that the concatenation of any such realized $M$ environments is itself a $(\beta,\lambda)$-Hölder environment over $T$ rounds.
@@
First, since $\varphi$ is $(\beta,\lambda/2)$-Hölder (assertion (a) of Tsybakov, 2009, Section 2.5), the reward function $t\mapsto\mu_{t,\textbf{a}}(a)$ for each arm $a$ is $(\beta,\lambda/2)$-Hölder being a sum of $(\beta,\lambda/2)$-Hölder functions with disjoint supports. Then, the gap functions $t\mapsto\max_{a^{\prime}}\mu_{t,\textbf{a}}(a^{\prime})-\mu_{t,\textbf{a}}(a)$ are $(\beta,\lambda)$-Hölder as a difference of two $(\beta,\lambda/2)$-Hölder functions.
@@
In what follows, we’ll make use of environments which use both formulas $\varphi_{a,\textbf{a},i}(x)$ and $\tilde{\varphi}_{a,\textbf{a},i}(x)$. All reward random variables $Y_{t}(a)$ will be Bernoulli’s and we’ll only specify the means $\mu_{t}(a)$.
@@
Now, consider a sub-environment $\mathcal{E}_{1}$ over $n$ rounds (which we’ll for ease momentarily parametrize via $[n]$) on which arm $1$ is optimal with:
@@
$R_{\mathcal{E}_{1}}(\pi,n)+R_{\mathcal{E}_{a}}(\pi,n)\geq\frac{n\Delta}{4}(\mathbb{P}_{\mathcal{E}_{1}}(N_{n}(1)\leq n/2)+\mathbb{P}_{\mathcal{E}_{a}}(N_{n}(1)>n/2))\geq\frac{n\Delta}{8}\exp\left(-\operatorname{KL}(\mathcal{P}_{1},\mathcal{P}_{a})\right),$ (9)
@@
$\operatorname{KL}\left(\operatorname{Ber}\left(\frac{1}{2}-\varphi\left(\frac{t}{T}\right)\right),\operatorname{Ber}\left(\frac{1}{2}+\varphi\left(\frac{t}{T}\right)\right)\right)$ $=\left(\frac{1}{2}+\Delta_{t}\right)\log\left(\frac{1/2+\Delta_{t}}{1/2-\Delta_{t}}\right)+\left(\frac{1}{2}-\Delta_{t}\right)\log\left(\frac{1/2-\Delta_{t}}{1/2+\Delta_{t}}\right).$
@@
where $\mathcal{H}_{t}$ is the filtration of history of observations and decisions up to round $t$. Now, for $i\in[T/M-1]$, there must exist an arm $a\neq 1$ whose conditional arm-pull count $\mathbb{E}\left[\sum_{t=i\cdot M+1}^{(i+1)\cdot M}\boldsymbol{1}\{\pi_{t}=a\}\mid\mathcal{H}_{i\cdot M}\right]\leq\frac{n}{K-1}$. Then, conditional on $\mathcal{H}_{i\cdot M}$, we can design environments $\mathcal{E}_{1}$ and $\mathcal{E}_{a}$ as before and lower bound the conditional regret by $\Omega(n\cdot\tilde{\lambda}\cdot(n/T)^{\beta})$.

### v2

@@ 6   Gap-Dependent Dynamic Regret Bounds
where $L$ is the number of stationary phases, the $\ell$-th of which has gap profile $\{\delta_{\ell}(a)\}_{a\in[K]}$. Unfortunately, it is long been known in the switching bandit literature that, in the worst case, Equation 5 cannot be attained simultaneously for different values of $L$ (Garivier and Moulines, 2011; Lattimore and Szepesvári, 2020). The reason for such a hardness is intuitively because of the additional exploration required to detect unknown changes, which forces $\sqrt{T}$ regret.
@@
In our next contribution, we give answers to both these questions in terms of the significant shift oracle, introduced in Section 4. Recalling such an oracle roughly plays arms until they incur significant regret Equation 1, and restarts at each significant shift, we will see that a careful regret analysis of this oracle gives rise to a faster rate than Equation 5 which is achievable adaptively in so-called safe environments. While a safe environment is already defined earlier in Section 5 as that under which no significant shift occurs, we give a more refined agent-based definition here in Definition 5.
@@ 6.1   Related Work on Gap-Dependent Regret
To start, we give an account of some works which aim to achieve the restarting oracle rate Equation 5 under structured changes.
@@
  • Seznec et al. (2020) achieve the rate Equation 5 in (restless) rotting bandits with decreasing rewards.
@@
  • Allesiardo et al. (2017) show gap-dependent regret bounds in non-stationary environments with a unique best arm, which is more restrictive than our notion of safe environment (Definition 5). Their rate is $\tilde{O}(K/\Delta)$ where $\Delta$ is the minimal average gap over time $\Delta:=\min_{a\neq a^{*}}T^{-1}\sum_{t=1}^{T}\delta_{t}(a)$, which is similar to our new gap-dependent rate Equation 7.
@@ 6.2   Refined Regret Analysis of the Significant Shift Oracle
Notation 4.
Let $\mathcal{H}_{t}$ be be the $\sigma$-algebra generated by random reward variables $\{Y_{s}(a)\}_{s\leq t,a\in[K]}$ and exogenous time-varying randomness $\{\pi_{s}\}_{s\leq t}$, as used by an algorithm $\pi$.
We will use $t_{1},\ldots,t_{K}$ to denote ordered stopping times with respect to the filtration $\{\mathcal{H}_{t}\}_{t\in[T]}$ and use $\mathcal{S}_{1},\cdots,\mathcal{S}_{T}$ to denote random subsets of $[K]$ which are adapted to this filtration.
@@
Let $t_{0}\doteq 1$ and let $\mathcal{S}_{1}\doteq[K]$. Then, we will recursively define $t_{i}$ and $\mathcal{S}_{t}$ for $t>t_{i-1}$ as follows: a stopping time $t_{i}>t_{i-1}$ is called an eviction time w.r.t. initial time $t_{i-1}$ if $\mathcal{S}_{t_{i}-1}=\mathcal{S}_{t_{i}-2}=\cdots=\mathcal{S}_{t_{i-1}}$ and
$\forall a\in\mathcal{S}_{t_{i}-1},[s_{1},s_{2}]\subseteq[1,t_{i}-1]:\sum_{s=s_{1}}^{s_{2}}\frac{\delta_{s}(a)}{|\mathcal{S}_{s}|}$ $\leq C_{3}\sqrt{\sum_{s=s_{1}}^{s_{2}}\frac{\log(T)}{|\mathcal{S}_{s}|}}.$ (6)
$(t_{1},t_{2},\ldots,t_{K})$ are a sequence of eviction times with safe armsets $\mathcal{S}_{1}\supseteq\mathcal{S}_{2}\supseteq\cdots\supseteq\mathcal{S}_{T}$.
@@
Plainly speaking, Equation 7 captures the regret of the significant shift oracle, or specifically an elimination procedure which tracks the safe armsets $\mathcal{S}_{t}$ and uniformly explores $\mathcal{S}_{t}$ at round $t$. Note that Equation 7 is a random quantity as the $t_{i},\mathcal{S}_{t}$ may depend on the random rewards and exogenous randomness of some algorithm $\pi$. Despite this randomness, we will next show that for any valid $t_{i},\mathcal{S}_{t}$ satisfying Definition 4, the rate of Equation 7 recovers Equation 5 and is achievable adaptively in safe environments (Definition 5).
6.3   Properties of New Gap-Dependent Regret Rate (Proofs in Supplement)
Proofs for this section are found in Appendix C.
@@ 6.3   Properties of New Gap-Dependent Regret Rate (Proofs in Supplement)
($\mathscr{R}(\{t_{i}\}_{i\in[K]},\{\mathcal{S}_{t}\}_{t\in[T]},\pi)\leq C_{3}^{2}\sum_{\ell=1}^{L}\sum_{a:\delta_{\ell}(a)>0}\frac{\log(T)}{\delta_{\ell}(a)}.$
We next show Equation 7 recovers the $\sqrt{KT}$ regret bound in safe environments, defined here.
@@
$\forall[s_{1},s_{2}]\subseteq[1,T]:\sum_{t=s_{1}}^{s_{2}}\frac{\delta_{t}({a^{\sharp}})}{|\mathcal{G}_{t}|}\leq C_{4}\sqrt{\sum_{t=s_{1}}^{s_{2}}\frac{\log(T)}{|\mathcal{G}_{t}|}}.$ (8)
@@
A significant shift cannot occur in a safe environment. Indeed, taking $\mathcal{G}_{t}\equiv[K]$ in Equation 8 gives us $\sum_{t=s_{1}}^{s_{2}}\delta_{t}({a^{\sharp}})<C_{4}\sqrt{K\cdot\log(T)\cdot(s_{2}-s_{1}+1)}$ which is a generalization of the reversal of Equation 1.
@@
($\mathscr{R}(\{t_{i}\}_{i\in[K]},\{\mathcal{S}_{t}\}_{t\in[T]},\pi)\leq C_{5}\sqrt{KT\log(T)}.$
@@ 6.4   Elimination Achieves Gap-Dependent Regret Rate in Safe Environments
As already hinted up to this point, we posit that a randomized elimination procedure (Algorithm 3) similar to Algorithm 2 in fact attains the rate of Equation 7 in safe environments. The major key difference in Algorithm 3 is the absence of the replay scheduling required for META and the use of refined elimination thresholds which tightly capture the variance of estimation.
@@
Going into detail, we estimate the relative gap $\delta_{t}(a^{\prime},a)$ via the unscaled estimates $\hat{\delta}_{t}(a^{\prime},a)/|\mathcal{A}_{t}|$ where we recall $\hat{\delta}_{t}(a^{\prime},a)$ of Equation 3 is the importance-weighted estimate and $\mathcal{A}_{t}$ is the active set of arms at round $t$.
@@ 4       Evict bad arms:
5       $\mathcal{A}_{t}\leftarrow\mathcal{A}_{t}\backslash\left\{a\in[K]:\text{$\exists$ round $t_{0}\leq t$: $\underset{a^{\prime}\in\mathcal{A}_{t}}{\max}\sum_{s=t_{0}}^{t}\frac{\hat{\delta}_{s}(a^{\prime},a)}{|\mathcal{A}_{s}|}>C_{6}\sqrt{\sum_{s=t_{0}}^{t}\frac{\log(T)}{|\mathcal{A}_{s}|}}$}\right\}$.
@@
Given any safe bandit environment over $T$ rounds, letting $\pi$ be Algorithm 3, we have w.p. at least $1-1/T^{2}$, for some eviction times and safe armsets $\{t_{i}\}_{i\in[K]},\{\mathcal{S}_{t}\}_{t\in[T]}$:
$\sum_{t=1}^{T}\delta_{t}(\pi_{t})\leq C_{5}\cdot\left(\log(T)+\mathscr{R}(\{t_{i}\}_{i\in[K]},\{\mathcal{S}_{t}\}_{t\in[T]},\pi)\right).$
@@
On the other hand, as mentioned earlier, it is known in switching bandits that the restarting oracle rate Equation 5 cannot be achieved adaptively for unknown $L$. However, this does not contradict our findings because the constructed hard environment (Lattimore and Szepesvári, 2020, e.g. Theorem 31.2) is not safe, violating Definition 5. Thus, similar to Suk and Kpotufe (2022), we find that the notion of significant shift (which decides the safeness of an environment) characterizes difficult non-stationarity in a new sense. So long as such a shift does not occur, we can attain the faster rate Equation 7.
@@ 6.5   Lower Bound for Gap-Dependent Regret Rate
We note the $\log(T)$ factor in Equation 6 of Definition 4 was only included for the sake of showing the regret upper bounds established up to this point. Going forward, we will ignore the $\log(T)$ factor when we refer to Equation 6.
@@
Let $\{t_{i}\}_{i=1}^{K}$ be an arbitrary set of rounds such that $t_{i+1}-t_{i}+1\geq K$ for all $i\in[K]$ with the convention that $t_{0}\doteq 1$ and $t_{K-1}=t_{K}\doteq T+1$. Fix a positive real number $R$ such that $R\leq\sum_{i=1}^{K-1}\sqrt{(t_{i}-t_{i-1})\cdot(K+1-i)}$. Let $\mathscr{E}$ be the class of environments such that (a) $t_{1},\ldots,t_{K}$ are valid deterministic eviction times with $C_{3}=(K-2)^{1/2}$ in Equation 6 for some shrinking sequence of safe armsets $\mathcal{S}_{1}\supseteq\cdots\supseteq\mathcal{S}_{T}$ and (b) such that:
@@ 7   Achievability of Gap-Dependent Rate in terms of Smoothness
Theorem 9 (proof in supplement).
@@
Acknowledgements
We thank Samory Kpotufe and anonymous referees for useful feedback on the manuscript. We also acknowledge computing resources from Columbia University’s Shared Research Computing Facility project, which is supported by NIH Research Facility Improvement Grant 1G20RR030893-01, and associated funds from the New York State Empire State Development, Division of Science Technology and Innovation (NYSTAR) Contract C090171, both awarded April 15, 2010.
@@ Appendix A Proof of Dynamic Regret Lower Bound (Theorem 1)
The construction will rely on bump reward functions which also appear in the classical minimax lower bounds for integrated risk of nonparametric regression of $(\beta,\lambda)$-Hölder functions (Tsybakov, 2009, e.g., Section 2.5). This will be combined with a Le Cam’s method style of argument for establishing a regret lower bound for stationary bandits (Lattimore and Szepesvári, 2020, Theorem 15.2), which we will modify to work for segments of mild non-stationarity.
@@
At a high level, we will construct $M$ instances of a randomly selected $(\beta,\lambda)$-Hölder environment of length $T/M$. we will then argue that the concatenation of any such realized $M$ environments is itself a $(\beta,\lambda)$-Hölder environment over $T$ rounds.
@@
First, since $\varphi$ is $(\beta,\lambda/2)$-Hölder (Tsybakov, 2009, Section 2.5, assertion (a)), the reward function $t\mapsto\mu_{t,\textbf{a}}(a)$ for each arm $a$ is $(\beta,\lambda/2)$-Hölder being a sum of $(\beta,\lambda/2)$-Hölder functions with disjoint supports. Then, the gap functions $t\mapsto\max_{a^{\prime}}\mu_{t,\textbf{a}}(a^{\prime})-\mu_{t,\textbf{a}}(a)$ are $(\beta,\lambda)$-Hölder as a difference of two $(\beta,\lambda/2)$-Hölder functions.
@@
In what follows, we will make use of environments which use both formulas $\varphi_{a,\textbf{a},i}(x)$ and $\tilde{\varphi}_{a,\textbf{a},i}(x)$. All reward random variables $Y_{t}(a)$ will be Bernoulli’s and we will only specify the means $\mu_{t}(a)$.
@@
Now, consider a sub-environment $\mathcal{E}_{1}$ over $n$ rounds (which we will for ease momentarily parametrize via $[n]$) on which arm $1$ is optimal with:
@@
$R_{\mathcal{E}_{1}}(\pi,n)+R_{\mathcal{E}_{a}}(\pi,n)$ $\geq\frac{n\Delta}{4}(\mathbb{P}_{\mathcal{E}_{1}}(N_{n}(1)\leq n/2)+\mathbb{P}_{\mathcal{E}_{a}}(N_{n}(1)>n/2))$
$\geq\frac{n\Delta}{8}\exp\left(-\operatorname{KL}(\mathcal{P}_{1},\mathcal{P}_{a})\right),$ (10)
@@
$\operatorname{KL}$ $\left(\operatorname{Ber}\left(\frac{1}{2}-\varphi\left(\frac{t}{T}\right)\right),\operatorname{Ber}\left(\frac{1}{2}+\varphi\left(\frac{t}{T}\right)\right)\right)$
$=\left(\frac{1}{2}+\Delta_{t}\right)\cdot\log\left(\frac{1/2+\Delta_{t}}{1/2-\Delta_{t}}\right)+\left(\frac{1}{2}-\Delta_{t}\right)\cdot\log\left(\frac{1/2-\Delta_{t}}{1/2+\Delta_{t}}\right).$
@@
where $\mathcal{H}_{t}$ is the filtration of history of observations and decisions up to round $t$. Now, for $i\in[T/M-1]$, there must exist an arm $a\neq 1$ whose conditional arm-pull count
$\mathbb{E}\left[\sum_{t=i\cdot M+1}^{(i+1)\cdot M}\boldsymbol{1}\{\pi_{t}=a\}\mid\mathcal{H}_{i\cdot M}\right]\leq\frac{n}{K-1}.$
Then, conditional on $\mathcal{H}_{i\cdot M}$, we can design environments $\mathcal{E}_{1}$ and $\mathcal{E}_{a}$ as before and lower bound the conditional regret by $\Omega(n\cdot\tilde{\lambda}\cdot(n/T)^{\beta})$.